# Reconnaissance with JWST of the J-region Asymptotic Giant Branch in Distance Ladder Galaxies: From Irregular Luminosity Functions to Approximation of the Hubble Constant

### 09 Jan 2024 v1

@@ 5.2 The JAGB in Perspective
We hypothesize that these variations in the shapes of the JAGB luminosity function may result from intrinsic, astrophysical differences in the distribution of carbon star demographics. Parada et al. (2021) suggested that some of the field-to-field variations in the measured JAGB luminosity seen in the LMC, listed in their Table 5 and described in their Fig. 5.1, may be due to variations in the star formation history in different regions. The differences are unlikely to result from interstellar reddening, due to the use of reddening maps to remove absorption. We would also not expect interstellar reddening to play a role in the host studied here, based on estimates of the variation in interstellar reddening with radial separation presented by Ménard et al. (2010) (see equation 29). For example, at semi-minor axes separations of 39 and 48 from the centers of NGC 1559 and NGC 5584 corresponding to the ellipses in Table 7, we expect NIR internal extinctions of 6 mmag and 5 mmag, respectively. Parada et al. (2021) also suggest that metallicity can drive the skewness of the J-region luminosity function and the JAGB luminosity, and choose either to use the LMC or SMC as anchors based on the skewness of the J-region luminosity functions in their target galaxies. We also note that it may also be possible that the J-region contains a mixed population of stars, such that not all groups of stars have the same mean magnitude. Further work is needed to investigate the origin of the variations in the measured JAGB and to demonstrate that these differences in treatment of different targets do not create undue variations and uncertainties in the estimated relative distances. Because we do not know whether the often asymmetric shape of the LF is due to the spread of a single population or the combination of multiple symmetric populations, it is not clear a priori which reference flux should provide the best standard candle, something that is crucial to take into consideration when using the JAGB as a standard candle.
@@
For an overall perspective, we plot in Fig. 15 the JAGB distances measured here against the Cepheid distances from Riess et al. (2022), adjusted for a NGC 4258 only anchor, together with the JAGB and Cepheid distances from Zgirski et al. (2021). For our distance measurements we use the weighted average of the distances shown in Fig. 11, and include the contribution of methodological variation in the estimated errors. We find that the JAGB and Cepheid distances show reasonable agreement within their error bars.
@@
The JWST data presented above demonstrate that until JAGB LF variations are standardized or understood, they should be included in the uncertainties when the JAGB is used to measure extragalactic distances and the Hubble constant. At present, this issue lay in the critical path to reach the few percent level precision in measuring the Hubble constant.
@@
Such issues are not unexpected for the development and refinement of a new standard candle. Recent work has already made inroads into improving the choice of radius at which to measure JAGB (Lee, 2023) (see discussion comparing radii in Appendix B) as well as how best to match the shape of the luminosity function to measure distances (Parada et al., 2023). Future studies and JAGB observations are likely to continue improving this promising standard candle.
Appendix A Effects of Smoothing on J-region Luminosity Functions
@@
We find that the choice of smoothing in the mode-based method described in Lee et al. (2022); Lee (2023) can bias JAGB measurements for asymmetric luminosity functions. We demonstrate this by plotting in Fig. 16 the smoothed luminosity function for NGC 4258, Inner+Outer West field in F150W using baseline color cuts of 1 mag and 1.5 mag, magnitude cuts of 21 mag and 24 mag, and three levels of smoothing ( 0.2, 0.3, and 0.4 mag). We note that smoothings of 0.2 mag is used by Lee (2023) and 0.25 and 0.3 mag are used by Lee et al. (2023). The difference between the brightest and faintest measured JAGBs for the range of smoothing from 0.2 and 0.4 mag is 0.09 mag. We caution that using different smoothing parameters with this measurement method across the distance ladder can potentially introduce biases if luminosity functions are not equally asymmetric in the anchor and host galaxies.
Appendix B Stability of the JAGB with Radius
@@
We perform a similar test with our sample by constructing several annuli centered on the galaxy and measure the JAGB using the stars inside each annulus. We use the same color and magnitude cuts as described in the main body of this paper. For the purpose of this exercise, we used the mode based method and smoothing from Lee et al. (2022); Lee (2023) to measure the JAGB.

### 29 Feb 2024 v2

@@ 5.2 The JAGB in Perspective
We hypothesize that these variations in the shapes of the JAGB luminosity function may result from intrinsic, astrophysical differences in the distribution of carbon star demographics. Parada et al. (2021) suggested that some of the field-to-field variations in the measured JAGB luminosity seen in the LMC, listed in their Table 5 and described in their Section 5.1, may be due to variations in the star formation history in different regions. The differences are unlikely to result from interstellar reddening, due to the use of reddening maps to remove absorption. We would also not expect interstellar reddening to play a role in the host studied here, based on estimates of the variation in interstellar reddening with radial separation presented by Ménard et al. (2010) (see equation 29). For example, at semi-minor axes separations of 39 and 48 from the centers of NGC 1559 and NGC 5584 corresponding to the ellipses in Table 7, we expect NIR internal extinctions of 6 mmag and 5 mmag, respectively. Parada et al. (2021) also suggest that metallicity can drive the skewness of the J-region luminosity function and the JAGB luminosity, and choose either to use the LMC or SMC as anchors based on the skewness of the J-region luminosity functions in their target galaxies. We also note that it may also be possible that the J-region contains a mixed population of stars, such that not all groups of stars have the same mean magnitude. Further work is needed to investigate the origin of the variations in the measured JAGB and to demonstrate that these differences in treatment of different targets do not create undue variations and uncertainties in the estimated relative distances. Because we do not know whether the often asymmetric shape of the LF is due to the spread of a single population or the combination of multiple symmetric populations, it is not clear a priori which reference flux should provide the best standard candle, something that is crucial to take into consideration when using the JAGB as a standard candle.
@@
For an overall perspective, we plot in Fig. 14 the JAGB distances measured here against the Cepheid distances from Riess et al. (2022), adjusted for a NGC 4258 only anchor, together with the JAGB and Cepheid distances from Zgirski et al. (2021). For our distance measurements we use the median of the distances shown in Fig. 10, and include the contribution of methodological variation in the estimated errors. We find that the JAGB and Cepheid distances show reasonable agreement within their error bars.
@@
The JWST data presented above demonstrate that until JAGB luminosity function variations are standardized or understood, they should be included in the uncertainties when the JAGB is used to measure extragalactic distances and the Hubble constant. At present, this issue lay in the critical path to reach the few percent level precision in measuring the Hubble constant.
@@
Such issues are not unexpected for the development and refinement of a new standard candle. Recent work has already made inroads into improving the choice of radius at which to measure JAGB (Lee, 2023) (see discussion comparing radii in Appendix C) as well as how best to match the shape of the luminosity function to measure distances (Parada et al., 2023). Future studies and JAGB observations are likely to continue improving this promising standard candle.
Appendix A Asymmetry and J-region Tilts
In this section, we investigate the shape and structure of the J-region in the CMD and its affect on its use as a standard candle. Ideally the shape would be simple (i.e., a fixed point) or if not, at least consistent between hosts.
We first quantify the color dependence of the JAGB reference magnitude in our data, showing in the upper row of Fig. 15 the JAGB reference magnitudes for three equal color bins in the range of 1 1.5 mag relative to their means. The JAGB reference magnitudes for the bins use the clipped mean and mode () as illustrative examples. For the five hosts we see a semi-quadratic relation between magnitude and color, with the center color bin being brighter than the mean by mag. This finding is consistent with the lack of any significant Pearson correlation between brightness and color (i.e., not a linear trend) shown in Fig. 5. We perform a rectification of the JAGB magnitudes based on color and a quadratic fit to these points (calculated separately for the clipped mean or mode) and plot the CMDs containing the rectified J-regions in Fig. 16. The Pearson correlation coefficients remain near zero after rectification.
We recalculate the median difference between the JAGB and Cepheid distances after color rectification in Fig. 15. Compared to without rectification, see Figure 11, we find little change. This is not surprising since the relation between color and luminosity is fairly consistent across hosts so the quadratic structure largely cancels when comparing measures over the same color range. We also find that removing this quadratic color structure does not change the asymmetry (most strongly present in the LF of NGC 4258), as evidenced by seeing the same 0.15 mag difference between the mode and mean.
For reference we repeat the same analysis for observations of the JAGB in the (1.15 microns) band, bluer but closer to the conventional -band (1.25 microns) than (1.50 microns). To do this we retrieved JWST GO-1995 observations of NGC 4536 and NGC 7250 (PI: W. Freedman Freedman et al., 2021) from MAST and perform photometry using DOLPHOT (Dolphin, 2000, 2016) previously studied by Lee et al. (2024). We use the same pipeline described in Section 2.2 and quality cuts and disk region selection from Lee et al. (2024) as applicable and measure the JAGB reference magnitude as a function of three color bins across = mag as studied by Lee et al. (2024). We show these measurements in Fig. 17. We find that the JAGB reference magnitudes in are not consistent with a flat line similar, to what we see with . The relative changes in the JAGB reference magnitudes with color in vs. are perhaps less consistent across galaxies than what we found with vs. in Fig. 15.
The non-constant magnitude of the JAGB with color demonstrated above for half a dozen galaxies in and leads us to revisit the expectation of a constant magnitude in the ground-based -band (1.25 microns) claimed in for example Madore & Freedman (2020). Weinberg & Nikolaev (2001) observed that the J-region in the K-band followed the relation:
where is a zero-point, and and refer to their respective bands. Madore & Freedman (2020) concluded from this relation that adopting a unit slope and rearranging Equation A1 yields J = , a constant. However, in detail this may not be broadly true for other hosts or in general.
We also note that the color dependence of the JAGB is not directly related to the asymmetry of some JAGB LFs, i.e., the former does not explain the latter. One can imagine an arbitrarily narrow color range in the center of the JAGB with a still asymmetric luminosity function. While we do not know the cause of the asymmetry, it may be caused by a mixture of different populations/types of carbon stars with different mean luminosities. Morgan et al. (2003) observed carbon stars in the LMC using the 2dF facility on the Anglo-Australian Telescope and found that the J-type carbon star sequence is offset by 0.6 mag from the K-type carbon star sequence on a K vs. J - K luminosity function. In addition, Fig. 10 in Ripoche et al. (2020) shows that both J-type and N-type carbon stars reside within the J-region used to measure the JAGB reference magnitude. Fig. 7 in Abia et al. (2022) shows that the bolometric luminosities of the J- and N-type carbon stars have different mean magnitudes. We retrieved the data from Abia et al. (2022) and find that this difference persists in the 2MASS J-band as well, with a difference in mean magnitudes between the J- and N-type carbon star luminosity functions mag at about 2 (see Fig. 18).
Finally, the spectral features distinctive to Carbon-rich stars are present in both the J, H, and bands suggesting that aby differences in the nature of the J-region studied above are not due to a lack of coverage of Carbon molecular bands. We refer the reader to the spectra from the X-shooter spectral library in Gonneau et al. (2016, 2017). For instance, in Fig. 5 in Gonneau et al. (2016) (seen in Fig. 19), we see that the representative carbon star in orange of group 3, which contains most of the J-region stars within the color range of 1.4 2 mag typically used to select J-region stars. Section 4.3 in Gonneau et al. (2016) notes that their group 3 stars, which have colors 1.6 mag and overlap with the colors used to define the J-region, have strong bands in both J- and H-bands suggesting that the spectral features of these Carbons stars do not preferentially lie in one of those two bands rather than the other. For context, the and bands cover between 1.013 to 1.282 m and 1.331 to 1.668 m, respectively.
Appendix B Effects of Smoothing on J-region Luminosity Functions
@@
We find that the choice of smoothing in the mode-based method described in Lee et al. (2022); Lee (2023) can bias JAGB measurements for asymmetric luminosity functions. We demonstrate this by plotting in Fig. 20 the smoothed luminosity function for the NGC 4258 Inner+Outer West field in F150W using baseline color cuts of 1 mag and 1.5 mag, magnitude cuts of 21 mag and 24 mag, and three levels of smoothing ( 0.2, 0.3, and 0.4 mag). For reference, we note that smoothings of 0.2 mag is used by Lee (2023) and 0.25 and 0.3 mag are used by Lee et al. (2024). The difference between the brightest and faintest measured JAGBs for the range of smoothing from 0.2 and 0.4 mag is 0.09 mag. We caution that using different smoothing parameters with this measurement method across the distance ladder can potentially introduce biases if luminosity functions are not equally asymmetric in the anchor and host galaxies.
Appendix C Stability of the JAGB with Radius
@@
We perform a similar test with our sample by constructing several annuli centered on the galaxy and measure the JAGB using the stars inside each annulus. We use the same color and magnitude cuts as described in the main body of this paper. For the purpose of this exercise, we used the mode based method and smoothing of ( = 0.25 mag) (Lee et al., 2024) to measure the JAGB.